Substitution Method
4m+n=11
m-8n=77

Answer by mananth(16949)

4 x + 1 y = 11 .............1
-8 y = 77 -1 x ---------2
/ -8 =
y = -9.63 + 0.13 x
Substitute y in (1)
4 x + 1 ( -9.63 + 0.13 x ) = 11
4 x -9.63 + 0.13 x = 11
4 x + 0.13 x = 11 9.63
4.13 x = 20.63
/ 4.13
x= 5
Plug the value of x in (1)
4 x + 1 y = 11
4 * 5 + 1 y = 11
20 + 1 y = 11
1 y = -9
/ 1
y= -9
m.ananth@hotmail.ca